OpenStudy (anonymous):

(t-5)^2=81 do I take 5 from both sides or add it?

OpenStudy (anonymous):

Neither

myininaya (myininaya):

to get rid of the square, you must square root both sides

OpenStudy (anonymous):

no what you do first is you hvae to square root both sides . only then you can add 5 to both sides. so t - 5 = - or + sqrt(81) so t - 5 = + or negative 9 so t - 5 = 9 and therefore t = 14 and t - 5 = -9 and therefore t = -4 so t = 14 and -4 are both valid answers

OpenStudy (anonymous):

I see how you got it now I guess with similar problems I thought I did the same thing to this one too. How would I do one with 1/2(x-1)^5=16

OpenStudy (anonymous):

1/2(x-1)=16^(1/5)

myininaya (myininaya):

first multiply by the reciprocal of 1/2 on both sides so you have grouping symbol by itself on one of the equation then you will have (x-1)^5=2*16

myininaya (myininaya):

(x-1)^5=32

myininaya (myininaya):

now since this is raised to the 5th power with take the 5th power of both sides giving us x-5=32^(1/5)

myininaya (myininaya):

x-5=2

myininaya (myininaya):

so x=2+5

myininaya (myininaya):

x=7

OpenStudy (anonymous):

how did you get 2

myininaya (myininaya):

what is 32^(1/5)

myininaya (myininaya):

since 32^(1/5)=[2*2*2*2*2]^(1/5)=2

OpenStudy (anonymous):

got it when I put in calculator

OpenStudy (anonymous):

is this right 3z^2 =48 z=16

myininaya (myininaya):

z^2=48/3=16 so z=plus or minus 4

myininaya (myininaya):

hey but what happens when you check x=7 in the question we did before this one

myininaya (myininaya):

did you catch my mistake above?

myininaya (myininaya):

it was suppose to be x-1=(32)^(1/5)

OpenStudy (anonymous):

No I didn't but I was checking it

myininaya (myininaya):

x-1=2

myininaya (myininaya):

x=1+2=3

OpenStudy (anonymous):

It is x-1+2=3 LOL

OpenStudy (anonymous):

so it isn't 7

OpenStudy (anonymous):

it is 3

myininaya (myininaya):

so you got it good! i have to go eat. yes it is 3

OpenStudy (anonymous):

thank you boy I could use your help all the time have a good supper